Investigators working on commericially cultivating shrimp are interested in determining the best conditions in which for them to grow. They are interested in knowing what impact water temperature (either 25 C, 30 C) and the salinity level (10 %, 25 % and 40 %) have on the growth of shrimp. Three replicates were used for each of the 6 possible treatment combinations. Post-larval shrimp were placed in each of the 18 tanks and fed the same amount of food. After one month the overall weight of all the shrimp in a tank was measured and compared to how much they weighed previously. The total sample standard deviation was found to be 15.87. The cell means in milligrams for the experiment are given below.

Temperature 25 C 30 C Row Mean
Salinity 10% 56.25 75.69 65.97
Salinity 25% 73.59 93.00 83.30
Salinity 40% 78.83 105.11 91.97
Column Mean 69.56 91.27

The investigators want to know whether there is an interaction between temperature and salinity. Which of the following wouldWhat is the grand mean? [1 pt(s)] Submit Answer Tries 0/3 Calculate the degrees of freedom Salinity Temperature Interaction EPlease give me correct and detailed answer!!!!
Answer #1

data ý the that solo given Temperature 20c 300 Row mean 65.97 Salinity 10% 56.25 75.69 83.30 93.00 Salinity asilo 73.59 Salin2 410) Salinity = 31 Temperatione = 21 = 1 Interation (3-0) (2-) =) 2x1 = 2 Errol = 3+2(3-1) = 6(2) 10 (Yejo 3 5) Calculating2 15,87 X 17 3 =d281-5673 2102.9958 2. SSE= H281,567 3 - Q120.9589 57.6126 SSE 57.6126 MSE- drenal 12 4.80105 MS LAB) 23.4963